Which of the following can be the sides of a right triangle?
 1.5 cm, 2cm, 2.5 cm.
In the case of right-angled triangles, identify the right angles

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let a = 1.5 cm, b = 2 cm, c = 2.5 cm
Let us assume the largest value is the hypotenuse side i.e. b = 2.5 cm.
Then, by Pythagoras theorem,
b^2=a^2+c^2
2.5^2= 1.5^2+ 2^2
6.25 = 2.25 + 4
6.25 = 6.25
The sum of square of two side of triangle is equal to the square of third side,
∴The given triangle is right-angled triangle.
Right angle lies on the opposite of the greater side 2.5 cm.

Thanks